OpenStudy (anonymous):

Solve the system: y=x^2+8x+9 x-y=-3

OpenStudy (anonymous):

y= x +3 x + 3 = x^2 + 8x + 9

OpenStudy (anonymous):

x^2 + 7x + 6=0 (x + 1)(x+6)=0 x= -1,-6 when we put that in y=x+3 we get y = 2,-3

OpenStudy (anonymous):

thats your solution

OpenStudy (anonymous):

Very nice :)

OpenStudy (anonymous):

you are right shark :)